There are 24 df.
Probability lesson represent outcomes 4th grade math?
That's not a question; indeed, it isn't even a sentence. If you're looking for a lesson to teach to 4th-graders, you might try an educators' site instead.
i think 6 out of 9 right? then you simplify to equal 1 out of 3.:)
z score = (400-500)/80=-1.25
probability under curve to right of -1.25 = .8944
A fair coin is tossed 3 times how many possible outcomes are there?
Each toss has 2 outcomes; so the number of outcomes for 3 tosses is 2*2*2 = 8
How many possible outcomes can you get by tossing 5 coins?
There are 25 or 32 possible outcomes can you get by tossing 5 coins.
the first seven Fibonacci numbers are 1,1,2,3,5,8,13. 13 is a Fibonacci number.
X = 50 => Z = (50 - 70)/12 = -20/12 = -1.33
So prob(X < 50) = Prob(Z < -1.33...) = 0.091
A card is drawn at random from a deck of cards Find the probability of getting the King of heart?
A deck of cards has 52 cards (not counting jokers which are optional). so if you draw one card at random you have one chance in 52 of getting a specific card. 1/52 = 1.923%
Of course it is.
The first draw always has 10 possibilities. The difference comes in the second draw.
The second draw has 10 possibilities if the first one was replaced, but only 9 possibilities
if the first one wasn't replaced.
If two events A and B are mutually exclusive then they are independent?
No, independence means they are not related. Mutually exclusive means they cannot occur at the same time.
We can certainly simplify that term. 6x - 3x = 3x. And -4y + y = -3y. So the simplified expression is 3x-3y. You could also call this 3(x-y). That's as far as we can go, since we have no way to evaluate x or y without further information.
What is the probability of getting 24 heads in a row when flipping a fair coin?
1 in 224, which is 1 in 16777216.
What is the probability of rolling a 3 and flipping a head?
probability of rolling a 3 = 1/6
probability of flipping a head = 1/2
therefore, overall = 1/12
12/27 reduces to 4/9
numerator = the number of acceptable outcomes
denominator = total number of outcomes (12 boys + 15 girls)
If the probability of an event is 3 to 7 what are the odds it will not occur?
four out of seven times
What are the boundaries of 25.6 ounces?
25.55 and 25.65. Because a number of ounces is normally rounded and therefore, the value of the actual number of ounces could be anywhere between 25.55 up to (but not including) 25.65.